# Thinking twice inside the box: is Wigner’s friend really quantum?

### v1

@@ Abstract
There has been a surge of recent interest in the Wigner’s friend paradox, sparking several novel thought experiments and no-go theorems. The main narrative has been that Wigner’s friend highlights a counterintuitive feature that is unique to quantum theory, and which is closely related to the quantum measurement problem. Here, we challenge this view. We argue that the gist of the Wigner’s friend paradox can be reproduced without assuming quantum physics, and that it underlies a much broader class of enigmas in the foundations of physics and philosophy. To show this, we first consider several recently proposed extended Wigner’s friend scenarios, and demonstrate that their implications for the absoluteness of observations can be reproduced by classical thought experiments that involve the duplication of agents. Crucially, some of these classical scenarios are technologically much easier to implement than their quantum counterparts. Then, we argue that the essential structural ingredient of all these scenarios is a feature that we call “Restriction A”: essentially, that a physical theory cannot give us a probabilistic description of the observations of all agents. Finally, we argue that this difficulty is at the core of other puzzles in the foundations of physics and philosophy, and demonstrate this explicitly for cosmology’s Boltzmann brain problem. Our analysis suggests that Wigner’s friend should be studied in a larger context, addressing a frontier of human knowledge that exceeds the boundaries of quantum physics: to obtain reliable predictions for experiments in which these predictions can be privately but not intersubjectively verified.
@@ I Introduction
In 1961, Eugene Wigner introduced his now famous thought experiment [1], illustrating an important subtlety of what is usually called the quantum measurement problem [2, 3]: a friend (F) observes the outcome of a measurement on a quantum system S, either seeing a flash or not. How should a superobserver (say, Wigner) describe the situation? On the one hand, both S and F are quantum systems, and then it “[…] follows from the linear nature of the quantum mechanical equations of motion that the state of the object plus observer is, after the interaction, […]” in an entangled state of FS [1]. On the other hand, regarding the friend F, “the question whether he did or did not see the flash was already decided in his mind, before I asked him.” [1]. Thus, it seems as if the correct description of the quantum state after F’s measurement would be an updated quantum state that contains only a single term and not a superposition of alternatives. Indeed, as proposed by Deutsch [4] (see also [3]), the friend might even communicate with Wigner and send him a message like “I have seen a definite outcome”, which does not alter any subsequent statistical predictions unless he communicates which outcome he has seen. However, Wigner may decide to perform an entangled measurement on the total system FS to confirm the assignment of an entangled quantum state.
@@
Recently, there has been a resurgence of interest in the Wigner’s friend thought experiment and its potential implications. In 2016, Frauchiger and Renner [5] introduced a version involving four agents, showing that it is in general inconsistent for such agents to reason indirectly by pooling each others’ predictions, even if those predictions involve only statements of probability zero or one. Brukner [6] analysed a further multi-agent version of Wigner’s friend by combining Wigner’s setup with a Bell scenario, describing his result as a “no-go theorem for observer-independent facts”. Building on this work, Bong et al. [7] derive a similar conclusion based solely on actually observed events, thus demarcating the captured phenomenon from that of Kochen-Specker contextuality [8]. Many further recent publications have considered aspects of the Wigner’s friend scenario and potential resolutions of its apparently paradoxical predictions and interpretations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
It seems self-evident and is typically understood that Wigner’s friend is an enigma specific to quantum theory: after all, the thought experiments mentioned above all rely on characteristic quantum phenomena like superposition or entanglement or the violation of Bell-type inequalities. However, it is important to note that these quantum ingredients are not themselves at the core of interest of the thought experiments, but rather their consequences for the involved agents and their observations, and how the agents’ observations relate to each other and to the rest of the world. In particular, Bong et al. [7] show that the violation of a “Local Friendliness” inequality (as predicted to be observed by quantum physics) implies that three metaphysical assumptions cannot all be jointly true, including an assumption of “Absoluteness of Observed Events”. In this paper, we ask whether these consequences are specific to quantum theory, regardless of the fact that the phenomena that are used to prove them within quantum theory are specifically quantum. To do so, we will study several thought experiments that are not intrinsically quantum, but that posit similar microscopic interventions over agents: the classical duplication of agents (in several versions), and cosmology’s Boltzmann brain problem. We show that important structural features of the recent Wigner’s friend thought experiments are reproduced by, or are relevant for, these scenarios. More concretely, we argue that there is a common structural core to all these thought experiments and others: a feature that we call “Restriction A”. In a nutshell, Restriction A says that our physical theories do not, or cannot, always give us a probabilistic description of the observations of all agents. Conditional on assuming the validity of the other two metaphysical statements, the results by Bong et al. imply a violation of Absoluteness of Observed Events, which we argue is a (particularly dramatic) instance of Restriction A. However, we argue that there are also important examples of Restriction A beyond quantum physics, and, in particular, in classical scenarios.
@@
Before summarising the conclusions that we draw (and do not draw) from this, let us clarify the motivation for this work further with an example. Consider the following analogy between our notion of Restriction A and the concept of correlation. In many popular-scientific accounts of quantum theory, the notion of entanglement is explained in an overly simplified and hence incorrect way, similarly as follows: Given the two electrons in a singlet state, finding that one electron has spin-up allows us to infer instantaneously that the other electron has spin-down — a very puzzling and counterintuitive feature of “spooky action at a distance”. Here is how it follows from the mathematics of the state vector:… However, the feature that is described here is not entanglement, but correlation. One can certainly use quantum theory, and its prediction of entanglement, to derive that the phenomenon of correlation appears in physics, but by no means is entanglement necessary to obtain physical situations where correlation applies. In particular, the phenomenon just described can be obtained with a pair of shoes that are randomly packaged into two boxes and sent to two agents. Correlation is a phenomenon of probability theory, and therefore of immense importance in classical physics and everyday life. What is specifically quantum is not the phenomenon of correlation, but the specific form of correlations that quantum theory admits (namely, ones that violate Bell-type inequalities). Similarly, we argue here that the difficulty of describing all agents with a single joint probability distribution is not specific to quantum theory, even though some of its implementation details may be (such as for EWF arguments).
That is, while the violation of a Local Friendliness inequality can be applied to demonstrate an instance of Restriction A, we show that this feature appears in classical physics too. Note, it is not at all our goal to “explain” or to “demystify” Wigner’s friend scenarios, nor is it to make an ontological claim regarding what actually happens in the world during their implementation. Rather, we make a broader, structural claim: essentially a mathematical statement about the impossibility in some scenarios of defining an absolute probability distribution – both for the well-studied quantum setting, as well as for the lesser acknowledged classical domain. In particular, here we argue that this structural notion that we call Restriction A is also at the core of cosmology’s Boltzmann brain problem, and that it relates to several other topics in philosophy, such as self-locating beliefs or accounts of personal identity. These structural similarities suggest a research strategy that would complement the existing quantum-theory-centred work on Wigner’s friend: regard it as a special case of a more general enigma which can and should be addressed in a unified way.
@@
Note that there has been previous research on the question of whether certain aspects of WF thought experiments are specifically quantum. Lostaglio and Bowles [20] have shown that the original WF thought experiment, involving one Wigner and one Friend, admits a simple classical explanation: Wigner and Friend may simply be two agents with differing knowledge about the same underlying physical configuration. This scenario is widespread and far from mysterious, in particular in classical statistical physics. Furthermore, Hausmann et al. [21] have shown that in classical theories such as Spekkens’ toy model [22], multi-agent paradoxes like Frauchiger and Renner’s [5] cannot be reproduced, whereas more general theories such as boxworld, featuring beyond-quantum Bell nonlocality, admit even stronger forms of such paradoxes [23]. While these results are important contributions to the WF research program, they do not meet the goal that we are setting ourselves in this paper: they study whether the corresponding theory admits the statistical prerequisites that are typically used to derive WF phenomenology (essentially, statistical incompatibilities across different contexts, as in the violation of Bell inequalities), but they do not study directly whether those theories allow us to draw similar metaphysical conclusions that the extended WF thought experiments imply.
Our article is organised as follows. In Subsection II.1, we review the recent Local Friendliness research program, as one of the most compelling of the EWF-type arguments. In particular, we summarise the no-go theorem for quantum physics in terms of the three theory-independent assumptions of [7], and the four metaphysical assumptions of [24]. We then present some analogous thought experiments in Subsection II.2, involving the duplication of agents, and argue that some of the LF assumptions are already untenable even in a classical setting. More precisely, the absoluteness of thoughts cannot naively be assumed in such scenarios, for which the traditional notion of agents and personal identity do not apply. Next, in Subsection II.3, we extend this argumentation and present a modified Sleeping Beauty thought experiment, where we argue similarly that core assumptions concerning the consistency of agents’ predictions are undermined classically (c.f. the Frauchiger-Renner Gedankenexperiment). In Subsection II.4, we argue why we ought to take these thought experiments seriously, alongside their quantum counterparts, and in Subsection II.5, we discuss further the relevance of considering these analogies at all. Next, in Subsection III.1, we offer a new perspective on the common structural properties of the quantum and classical thought experiments, leading us in particular to the definition of a structural property of physical theories which we call “Restriction A”. We discuss Wigner’s friend scenarios in these new terms in III.2, and demonstrate the relevance of these restrictions beyond quantum theory in Subsection III.3, focusing on the example of cosmology’s Boltzmann brain problem. Finally, we conclude in Section IV.
@@ II Wigner’s friends: quantum and classical
In this section, after reviewing the EWF scenario of [7, 24], we describe several classical thought experiments that we argue reproduce main implications of the extended quantum Wigner’s friend experiments. A thorough structural analysis (in terms of what we call “Restriction A”) will follow in the subsequent Section III. For a comparison to existing literature (in particular, to Kent’s work [25]), see Appendix V.1.
@@ II.1 The quantum thought experiments
It is shown in [24] that the LF no-go theorem can also be expressed as an incompatibility between quantum theory and the conjunction of the four metaphysical assumptions. In particular, Ego Absolutism states that “my” (in the sense of the first person) communicable thoughts are absolutely real – i.e. my thoughts are objective and need not be qualified relative to anything. Meanwhile, Friendliness states that the communicated thoughts of other intelligent parties are equally as real as my own communicable thoughts. The two together imply that both Wigner’s and his friend’s thoughts (which will also contain correlates of their observations) should be taken as absolutely real. In conjunction with Physical Supervenience (that thoughts supervene on physical processes in a bounded region of spacetime), this gives us something metaphysically analogous to AOE. Therefore, when we also assume Local Agency, the contradiction with the predictions of quantum theory (c.f. [7]) can be recovered, this time as a “thoughtful” LF no-go theorem.
@@ II.2 Some classical thought experimentation
The contradiction presents an important challenge to interpretations of quantum theory, asking which of the six assumptions (four metaphysical, plus two technological) of [24] it is prepared to drop. We would like to make a case though for how similar metaphysical dilemmas arise classically too, simply by considering thought experiments in which persons “branch”. Our claim is that the notion of “my” (in my communicable thoughts) can be ambiguous, and that this may be one reason for the failure of the conjunction of the four metaphysical assumptions, quantumly but also classically. That is, one does not need to go to the quantum regime in order to see that the language with which we discuss persons and thoughts is inherently restricted, and runs us into contradictions when taken to more exotic scenarios.
@@
Naively understood, there exists an ambiguity regarding what “my” indexes for branching scenarios. Returning to the example of Freya, who is yet to be duplicated, she reads Ego Absolutism to say that her communicable thoughts are absolutely real. This includes her thoughts in that instance, such as “I am hungry”. It may also be understood to include thoughts she had this morning, such as “It is raining”. Does it include her future thoughts though? This afternoon, she will be duplicated, whereupon her future copies will have separate experiences. Thus, in describing any future thought she may have, there is an inherent ambiguity as to the meaning of such statements, and whether or not we should take their referent as “absolute”. That is, it is unclear what the words “my (future) thoughts”, if uttered by Freya before the start of the experiment, would refer to, and in disregarding this indexical ambiguity,we will typically be led to mathematical formulations of Ego Absolutism that tacitly involve additional assumptions. In particular, it will lead to the formal assumption that there is always, at every time, a single variable describing a single thought of some person called Freya, while in this branching scenario there are actually two. Indeed, this assumption is part of the mathematical formulation of Bong et al. [7].
Furthermore, notice the ambiguity of the referent “they” in the following principle:
Friendliness: [24] If, by open-ended communication, an independent party displays cognitive ability at least on par with my own, then they have thoughts, and any thought they communicate is as real as any communicable thought of my own.
Together with the previous assumption, we find a similar ambiguity here, with respect to other agents. A second agent, say Wigner, who interacts with Freya and deems her to exhibit sufficient cognitive ability, takes the conjunction of Ego Absolutism and Friendliness to say that Freya’s thoughts are absolutely real. Yet, statements pertaining to “Freya” likewise presuppose additional commitments regarding their referent – in particular that there exists a single variable that describes any thought “she” may have. This breaks down when we consider the type of drastic interventions over Freya that we posit for both quantum and classical scenarios.
To be more concrete, we now consider a simple (albeit difficult to implement) thought experiment within classical physics, which illustrates the described ambiguity:
[…]
Figure 2: Sketch of Thought experiment 1; inside the lab, Freya interacts with some system, causing her to be duplicated into two identical copies. Wigner asks her which side of the room she finds herself in, whereupon the slower copy to respond is destroyed, leaving just one copy of Freya left in the lab.
of absoluteness for this scenario, without an account of personal identity that elucidates a real, singular, persisting referent.
We can adopt strategies from philosophers, as discussed in Appendix V.2, in order to make sense of the situation. For example, by Parfit’s anti-realist account of personal identity [29], Wigner’s statement may be interpreted along the lines of “The Freya who matters to me was absolutely on the right side of the room”. Alternatively, Lewis [31] (whose ontology of persons resembles four-dimensional, potentially overlapping spacetime “worms”) might interpret Wigner’s claim to say something like “The Freya who informed me that she was on the right side of the room was absolutely on the right of the room”. Or Sider [32] (with a person ontology composed of three-dimensional “person-stages”) may instead argue “The Freya who existed instantaneously at the completion of the experiment was absolutely on the right side of the room”. However, a statement that just refers to “Freya” without a specific account of identity, is ambiguous in its referent, and provides an incomplete description of the experiment. Moreover, the possibility for two Freyas may even constitute a loophole in the LF no-go theorem, as suggested by Kent [25].
@@
Let us argue slightly more formally. Consider the physicist Wigner who believes that Ego Absolutism is true, and who observes that Freya is being prepared for an experiment (which will turn out to implement Thought Experiment 1). Wigner, however, is not aware of all the details of the experiment; in particular, the closed laboratory is to some extent a “black box” for him, so that he does not have a complete description of what goes on in it. In particular, he may not know about the duplication process. Then, formally, Wigner will think that there is a variable (perhaps called $c$) which will describe Freya’s thoughts (and observations) during the experiment. His subjective uncertainty about the outcome will lead Wigner to treat this as random variable, which will take either one value or another, $c=c_{L}$ or $c=c_{R}$, because this is what variables do in probability theory and hence in classical physics. Moreover, the value of this variable is absolute in the sense that we can think of it as belonging to a larger set of facts of the physical world, described e.g. by a collection of random variables that have a joint distribution in Kolmogorov’s sense. However, this description is not valid in Thought Experiment 1: there is no random variable $c$ of this kind. The particular intervention that occurs in the black box disrupts the usual notion of probability space that Wigner assumes, rendering statements like “the probability that $c=c_{R}$ is $p$” undefined, because $c$ becomes undefined. Following Kent [25], we might describe it as there being two random variables $c_{1}$ and $c_{2}$ instead, describing the thoughts of the two Freyas.
Next, let us consider another, similar thought experiment, now including “merging” (c.f. [28]) to explore an analogous operation to the unitary reversal posited by LF experiments. This time, we consider just Freya and her thoughts and beliefs around an experiment.
[…]

### v2

# On the significance of Wigner’s Friend in contexts beyond quantum foundations

@@ Abstract
There has been a surge of recent interest in the Wigner’s Friend paradox, sparking several novel thought experiments and no-go theorems. The main narrative has been that Wigner’s Friend highlights a counterintuitive feature that is unique to quantum theory, and which is closely related to the quantum measurement problem. Here, we challenge this view. We argue that the gist of the Wigner’s Friend paradox can be reproduced without assuming quantum physics, and that it underlies a much broader class of enigmas in the foundations of physics and philosophy. To show this, we first consider several recently proposed Extended Wigner’s Friend scenarios, and demonstrate that some of their implications for the absoluteness of observations can be reproduced by classical thought experiments that involve the duplication of agents. Crucially, some of these classical scenarios are technologically much easier to implement than their quantum counterparts. Then, we argue that the essential structural ingredient of all these scenarios is a feature that we call “Restriction A”: essentially, that a physical theory cannot give us a probabilistic description of the observations of all agents. Finally, we argue that this difficulty is at the core of other puzzles in the foundations of physics and philosophy, and demonstrate this explicitly for cosmology’s Boltzmann brain problem. Our analysis suggests that Wigner’s Friend should be studied in a larger context, addressing a frontier of human knowledge beyond quantum foundations: to obtain reliable predictions for experiments in which these predictions can be privately but not intersubjectively verified.
@@ I Introduction
In 1961, Eugene Wigner introduced his now famous thought experiment [1], illustrating an important subtlety of what is usually called the quantum measurement problem [2, 3]. A friend (F) observes the outcome of a measurement on a quantum system S, either seeing a flash or not. How should a superobserver (say, Wigner) describe the situation? On the one hand, both S and F are quantum systems, and then it follows from linearity that the system and the Friend should be described by the entangled state FS. On the other hand, regarding the Friend, “the question whether he did or did not see the flash was already decided in his mind, before [Wigner] asked him” [1]. Thus, it seems as if the correct description of the quantum state after F’s measurement would be an updated quantum state that contains only a single term and not a superposition of alternatives. Indeed, as proposed by Deutsch [4] (see also [3]), the Friend might even communicate with Wigner and send him a message such as “I have seen a definite outcome”. Providing the Friend does not communicate which outcome they have seen, this does not alter any subsequent statistical predictions. Therefore this information does not change Wigner’s state assignment, yet at the same time challenges his description of an indefinite Friend.
@@
Recently, there has been a resurgence of interest in the Wigner’s Friend (WF) thought experiment and its potential implications, via so-called Extended Wigner’s Friend (EWF) scenarios. In 2016, Frauchiger and Renner [5] introduced a version involving four agents, showing that it is in general inconsistent for such agents to reason indirectly by pooling each others’ predictions, even if those predictions involve only statements of probability zero or one. Brukner [6] analysed a further multi-agent version of Wigner’s Friend by combining Wigner’s setup with a Bell scenario, describing his result as a “no-go theorem for observer-independent facts”. Building on this work, Bong et al. [7] derive a similar conclusion based solely on actually observed events, thus demarcating the captured phenomenon from that of Kochen-Specker contextuality [8]. Many further recent publications have considered aspects of the Wigner’s Friend scenario and potential resolutions of its apparently paradoxical predictions and interpretations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 20, 19].
@@
It seems self-evident and is typically understood that Wigner’s Friend is an enigma specific to quantum theory: after all, the thought experiments mentioned above all rely on characteristic quantum phenomena, such as superposition or entanglement, or the violation of Bell-type inequalities. However, in this paper we ask whether the consequences for the involved agents and their observations could also be achieved in different settings, without quantum theory. In particular, a substantial part of the community take EWF arguments to indicate that quantum theory is incompatible with certain notions of “absoluteness” of observations, and of how the observations of different agents relate to one another. We ask though whether this is a uniquely quantum feature, or whether it could be true of other physical theories – even of classical probability theory. To do so, we will study several thought experiments that are not intrinsically quantum, but that posit similar microscopic interventions over agents: the classical duplication of agents (in several versions), and cosmology’s Boltzmann brain problem. We show that important structural features of EWF experiments are reproduced by these scenarios.
More concretely, we argue that there is a common structural core to all these thought experiments and others: a feature that we call “Restriction 1”. In a nutshell, Restriction 1 says that our physical theories cannot always give us a probabilistic description of the observations of all agents. Conditional on assuming the validity of the other two metaphysical statements, the results by Bong et al. imply a violation of Absoluteness of Observed Events, which we argue is a (particularly dramatic) instance of Restriction 1. However, we argue that there are also important examples of Restriction 1 beyond quantum physics, and, in particular, in classical scenarios.
@@
Before summarising our conclusions, let us clarify the motivation for this work further with an example. Consider the following analogy between our notion of Restriction 1 and the concept of correlation. In many popular-scientific accounts of quantum theory, the notion of entanglement is explained in an overly simplified and hence incorrect way, similarly as follows: Given the two electrons in a singlet state, finding that one electron has spin-up allows us to infer instantaneously that the other electron has spin-down — a very puzzling and counterintuitive feature of “spooky action at a distance”. Here is how it follows from the mathematics of the state vector:… However, the feature that is described here is not entanglement, but correlation. One can certainly use quantum theory, and its prediction of entanglement, to derive that the phenomenon of correlation appears in physics, but by no means is entanglement necessary to obtain physical situations where correlation applies. In particular, the phenomenon just described can be obtained with a pair of shoes that are randomly packaged into two boxes and sent to two agents. Correlation is a phenomenon of probability theory, and therefore of immense importance in classical physics and everyday life. What is specifically quantum is not the phenomenon of correlation, but the specific form of correlations that quantum theory admits (namely, ones that violate Bell-type inequalities). Similarly, we argue here that the difficulty of describing all agents with a single joint probability distribution is not specific to quantum theory, even though some of its implementation details are (such as the violation of Local Friendliness inequalities). That is, while quantum EWF scenarios demonstrate (under further assumptions) instances of Restriction 1, we show that this feature appears in classical physics too.
Our paper is divided into two main sections:
  • In Section II, we look closely at the metaphysical assumptions of EWF scenarios. We argue that some of the assumptions for EWF scenarios, regarding the absoluteness of observations (Subsection II.2) or the consistency of agents’ beliefs (Subsection II.3), also fail to hold in classical thought experiments. In particular, we present two thought experiments, which involve the classical duplication of agents. We argue that, due to the absence of the traditional notion of personal identity, the assumptions are untenable even without any quantum ingredients.
  • In Section III, we formulate a new, structural property of physical theories, which we call Restriction 1. This is a mathematical statement about the impossibility of defining a probability distribution for the observations of agents in certain situations. We argue that Restriction 1 is the structural core of EWF scenarios (notably including Local Friendliness and the Frauchiger-Renner thought experiment), as well as of other, classical puzzles in physics and philosophy (such as the Boltzmann brain problem).
  • From this, we conclude (see Section IV) that Wigner’s Friend has significance far beyond quantum physics. EWF scenarios demonstrate in a particularly dramatic way that our current physical theories cannot always tell agents what they should expect to observe (Restriction 1 for $N=1$ observer). Moreover, even when they do, those predictions cannot be obtained as marginals of a joint prediction for the observations of all agents (Restriction 1 for $N\geq 2$ agents). We conclude that Wigner’s Friend should be studied alongside other physics and philosophy puzzles that feature Restriction 1, to mutual benefit.
@@
Note that there has been previous research on the question of whether certain aspects of WF thought experiments are specifically quantum. Lostaglio and Bowles [21] have shown that the original WF thought experiment, involving one Wigner and one Friend, admits a simple classical explanation: Wigner and Friend may simply be two agents with differing knowledge about the same underlying physical configuration. This scenario is widespread and far from mysterious, in particular in classical statistical physics. Furthermore, Hausmann et al. [22] have shown that in classical theories such as Spekkens’ toy model [23], multi-agent paradoxes like Frauchiger and Renner’s [5] cannot be reproduced, whereas more general theories such as boxworld, featuring beyond-quantum Bell nonlocality, admit even stronger forms of such paradoxes [24]. While these results are important contributions to the WF research program, they do not meet the goal that we are setting ourselves in this paper: they study whether the corresponding theory admits the statistical prerequisites that are typically used to derive WF phenomenology (essentially, statistical incompatibilities across different contexts, as in the violation of Bell inequalities), but they do not study directly whether those theories allow us to draw similar metaphysical or structural conclusions that the Extended WF thought experiments imply. Neverthless, our research is closely related to the “toy model” program of Spekkens and coauthors, in that we ask which aspects of the EWF arguments can be reproduced classically, versus which cannot – with the caveat that here we study features of theories, rather than phenomena. We comment more on this in Subsection II.5.
@@ II Wigner’s friends: quantum and classical
In this section, after reviewing the EWF scenario of [7, 25] in Subsection II.1, we describe several classical thought experiments that we argue reproduce main implications of the extended quantum Wigner’s Friend experiments (Subsections II.2 and II.3). In Subsection II.4, we argue why these thought experiments ought to be taken seriously alongside their quantum counterparts, and, in Subsection II.5, discuss the relevance of considering these classical analogues in the first place. A thorough structural analysis (in terms of what we call “Restriction 1”) will follow in the subsequent Section III. For a comparison to existing literature that already relates quantum phenomena with duplication (in particular, Kent’s work [26]), see Appendix V.1.
@@ II.1 The quantum thought experiments
Moreover, they propose the following additional technological assumptions for its implementation:
  1. 5.
    Human-level AI can be practically implemented on a digital computer.
  2. 6.
    Universal quantum computing is physically possible at very large scale, and very fast.
It is shown in [25] that the LF no-go theorem can also be expressed as an incompatibility between quantum theory and the conjunction of the four metaphysical assumptions. In particular, Ego Absolutism states that “my” (in the sense of the first person) communicable thoughts are absolutely real – i.e. my thoughts are objective and need not be qualified relative to anything. Meanwhile, Friendliness states that the communicated thoughts of other intelligent parties are equally as real as my own communicable thoughts. The two together imply that both Wigner’s and his Friend’s thoughts (which will also contain correlates of their observations) should be taken as absolutely real. In conjunction with Physical Supervenience (that thoughts supervene on physical processes in a bounded region of spacetime), this gives us something metaphysically analogous to AOE. Therefore, when we also assume Local Agency, the contradiction with the predictions of quantum theory (c.f. [7]) can be recovered. Together with the two technological assumptions regarding its implementation, this gives us the “thoughtful” LF no-go theorem: one cannot consistently hold all 6 assumptions.
@@ II.2 Some classical thought experimentation
The contradiction presents an important challenge to interpretations of quantum theory, asking which of the six assumptions of [25] it is prepared to drop. We would like to make a case though for how similar metaphysical dilemmas arise classically too, simply by considering thought experiments in which persons “branch”. Our claim is that the notion of “my” (in my communicable thoughts) can be ambiguous, and that this may be one reason for the failure of the conjunction of the four metaphysical assumptions, quantumly but also classically. That is, one does not need to go to the quantum regime in order to see that the language with which we discuss persons and thoughts is inherently restricted, and runs us into contradictions when taken to more exotic scenarios.
@@
Figure 2: Sketch of Thought Experiment II.2; inside a lab, Freya is duplicated into two identical copies, who wake up in different coloured rooms. This procedure is then reversed, such that a singular Freya wakes up again in a new lab room.
@@
Naively understood, there exists an ambiguity regarding what “my” indexes for branching scenarios. Returning to the example of Freya, who is yet to be duplicated, she reads Ego Absolutism to say that her communicable thoughts are absolutely real. This includes her thoughts in that instance, such as “I am hungry”. It may also be understood to include thoughts she had this morning, such as “It is raining”. Does it include her future thoughts though? This afternoon, she will be duplicated, whereupon her future copies will have separate experiences. Thus, in describing any future thought she may have, there is an inherent ambiguity as to the meaning of such statements, and whether or not we should take their referent as “absolute”. That is, it is unclear what the words “my (future) thoughts”, if uttered by Freya before the start of the experiment, would refer to, and in disregarding this indexical ambiguity, we will typically be led to mathematical formulations of Ego Absolutism that tacitly involve additional assumptions. Note too that the same ambiguity can be found in the Friendliness assumption, when “an independent [and intelligent] party” is the subject of branching. These claims lead to the formal assumption that there is always, at every time, a single variable describing a single thought of some person called Freya, while in this branching scenario there are actually two. Indeed, this assumption is part of the mathematical formulation of Bong et al. [7].
Let us now collect these ideas into the following simple, albeit difficult to implement, classical thought experiment involving both duplication and “merging” (c.f. [29]) :
{mdframed}
Thought Experiment 1. Consider a classical agent, Freya, who prepares to perform an experiment inside a closed laboratory. We refer to Freya at the start of the experiment as Freya0 in order to distinguish her from subsequent versions of herself. In her lab room, Freya0 is duplicated via some classical cloning device, following which one copy (FreyaB) wakes up in a blue room, and the other (FreyaG) in a green room. Upon awakening, each copy observes which colour room they find themselves in. Next, the two copies are put back to sleep and the memories of their respective experiences erased. They are then merged back together into a singular Freya, who wakes up again in the exit lab room, without any recollection of observations in either colour room in between.
The interventions in the experiment run us into problems concerning personal identity. Freya, reflecting on the experiment, would surely believe herself to be the same person that arrived at the lab to be duplicated, since she has sufficient physical and psychological continuity with Freya0. However, she has no recollection of waking in either blue or green rooms, therefore she cannot identify herself solely with either FreyaB or FreyaG. Still, “forgetting” is not generally thought to be sufficient to deem someone to be a different person. Moreover, in all other ways Freya has a very high degree of physical and psychological continuity with the two copies who were in the coloured rooms, and is causally continuous with them. In some sense, she is (/was) both FreyaB or FreyaG. The truth or falsity of statements such as “I was in the green room”, “I was in the blue room”, or even “I was in the blue and the green rooms” are ambiguous, without clarifying some specific account of personal identity that elucidates a real, singular, persisting referent. On such basis, simply looking at the wording of Ego Absolutism and Absoluteness of Observed Events, one may already regard these claims as being violated in the classical thought experiment.
More formally, of crucial importance are the structural and mathematical claims that are implicitly associated with the plain language statements: namely, that there is a single random variable at every moment of the experiment that describes Freya’s thoughts. This is explicit in Equation (1), and for the 4-party setup of [7]; AOE is taken to imply mathematically that there is a random variable $c$ (and $d$), which describes the thoughts of the friend(s) throughout the experiment. This assumption is manifestly violated in our classical thought experiment, despite looking the same from the outside as elements of WF or LF experiments.
@@
Let us consider an external perspective. Consider the physicist Wigner, outside Freya’s lab, who holds both Ego Absolutism and Friendliness. Wigner, however, is not aware of all the details of the experiment – only that there is a green and a blue room, one of which he presumes Freya must pass through to get to the exit room. To some extent though, the laboratory is a “black box” to Wigner, so that he does not have a complete description of what goes on inside it. Formally, Wigner might think that there is a variable (perhaps called $c$) that describes Freya’s observations during the experiment. His subjective uncertainty about the outcome would lead Wigner to treat this as random variable, taking either one value or another, $c=c_{B}$ or $c=c_{G}$, since this is what variables do in probability theory and hence in classical physics. Moreover, the value of this variable is absolute in the sense that we can think of it as belonging to a larger set of facts of the physical world, described e.g. by a collection of random variables that have a joint distribution in Kolmogorov’s sense. However, this description is not valid in Thought Experiment II.2: there is no random variable $c$ of this kind that Wigner could ascribe. The particular intervention that occurs in the black box disrupts the usual notion of probability space that Wigner might assume, hence rendering statements like “the probability that $c=c_{B}$ is $p$” undefined, because $c$ becomes undefined. Following Kent [26], we might describe it as there being two random variables $c_{1}$ and $c_{2}$ instead, describing the thoughts of the two Freyas.